OpenStudy (anonymous):

I HATE THIS, someone please help? Find an equivalent expression write your answer as a fraction. 9^-3

OpenStudy (amistre64):

- exponents go on the bottom of a fraction: 9^-3 means (1/9^3) = 1/(9.9.9) = 1/729

OpenStudy (id21):

\[(1/9)^{3}= 1/729\]

OpenStudy (anonymous):

Thank you :D